Use the quadratic formula to find all degree solutions and θ if

0° ≤ θ < 360°.

Use a calculator to approximate all answers to the nearest tenth of a degree. (Enter your answers as a comma-separated list. If there is no solution, enter NO SOLUTION.)

2 sin2 θ − 2 sin θ − 1 = 0
Answer #1

Given zsinho - 2 sin 0-1 =0 , 0ʻ<0. < 360° substitute u for sino 24-24-ico -biſ 6²-yac O za -(-2) + VE2)*–462)(-1) 2 (12) u=sino = I-B O= sin-(1-P) O- -21.5 The sine function is positive in second quadrants. To find the solution, subtrat the referen